Circular No. 6972

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998cy, 1998cz, 1998da, 1998db, 1998dc N. Regnault and O. Perdereau, Laboratoire de l’Accélérateur Linéaire,

Université de Paris-Sud, on behalf of the EROS collaboration (cf. IAUC 6762, 6904), report the discovery of five apparent supernovae discovered in the course of their automated supernova search with the 1-m Marly telescope (+ two 1-deg^{2} CCD cameras) at the European Southern Observatory:

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1998cy May 20 14^{h}02^{m}42.^{s}82 − 5^{o}45′05′′.7 19.5: near center
1998cz May 21 12 03 06.50 −12 37 41.5 18.5: near center
1998da July 1 0 27 13.76 −39 25 25.6 20.5: 2′′.4 W, 1′′.2 S
1998db July 1 22 55 54.64 −39 26 29.8 19.4: near center
1998dc July 5 22 20 16.10 −44 33 56.1 19.5: 24′′.6 W, 3′′.9 N

SN 1998cz appeared in a frame constructed from the difference between CCD frames taken on 1997 Feb. 9 and 1998 May 21, and it was confirmed on frames taken on May 26. Likewise, SNe 1998cy, 1998da, 1998db, and 1998dc were not present on CCD frames taken on 1997 Feb. 3, 1997 Oct. 5, 1997 Oct. 22, and 1996 Oct. 8, respectively, and they were confirmed on frames taken on 1998 May 26, July 2, July 2, and July 6.

XTE J2123–058 J. A. Tomsick, J. Kemp, and J. P. Halpern, Columbia University; and

D. Hurley-Keller, University of Michigan, report on the orbital period of this x-ray transient (IAUC 6955, 6958) from photometry of its optical counterpart (IAUC 6957): “We obtained V -band data at the MDM Observatory on June 30–July 4, and on the Kitt Peak 0.9-m telescope on July 15 and 16. Quasi-sinusoidal variations of amplitude 0.9 mag are seen. The best-fitting period is 5.957 ± 0.003 hr. Times of minimum light reported by Casares et al. (IAUC 6971) are consistent with this period, and enable us to eliminate all of the 1-day aliased candidates, including their proposed values, from consideration. The mean light curve thus obtained is not characterized by narrow eclipses, but is similar to the smooth modulation seen in many lowmass x-ray binary systems, with a flat top lasting 0.25 cycle and a broad triangular minimum lasting 0.75 cycle. An ephemeris for the time of minimum light is HJD 2451009.888(3) + 0.2482(1)E. At maximum light, V has remained at ∼ 16.8 since discovery. We also see evidence for three optical bursts of amplitude 0.3 mag or greater on June 30 and July 1.”

1998 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green